OpenStudy (anonymous):

Complete the conjecture: The product of two odd numbers is _____. A.odd B.prime C.even D.negative

OpenStudy (anonymous):

odd

OpenStudy (anonymous):

Do you know why?

OpenStudy (mertsj):

Multiply some odd numbers and figure it out.

OpenStudy (zzr0ck3r):

(2k+1) * (2J+1) when j and k are integers 4kj+2j+2k+1 = 2(2kj+k+j)+1 thus = 2z+1 for some z in integers... thus odd

OpenStudy (anonymous):

exactly as shown by zzr0ck3r

OpenStudy (zzr0ck3r):

note: any odd number can be written as 2k+1 where k is some integer.

OpenStudy (anonymous):

I DID AND IT'S ODD AS MY ANSWER

OpenStudy (zzr0ck3r):

:)